OpenStudy (chroncake):

maria has 20 coins in her pocket and all of her coins are either quarters or dimes. the coins total $3.20. how many of each coin does she have?

OpenStudy (rcp031):

Let x be the No. of quarters coins Let y be the No. of dimes. Then maria has 20 coins in her pocket means x+y=20 But, quarters means 1/25 of dollar and dime means 1/10 of a dollar Hence, x*(1/25)+y*(1/10)=3.20 Solve for x& y x+y=20 x*(1/25)+y*(1/10)=3.20